# From Pixel to Cancer: Cellular Automata in Computed Tomography

### v1

@@ Abstract
AI for cancer detection encounters the bottleneck of data scarcity, annotation difficulty, and low prevalence of early tumors. Tumor synthesis seeks to create artificial tumors in medical images, which can greatly diversify the data and annotations for AI training. However, current tumor synthesis approaches are not applicable across different organs due to their need for specific expertise and design. This paper establishes a set of generic rules to simulate tumor development. Each cell (pixel) is initially assigned a state between zero and ten to represent the tumor population, and a tumor can be developed based on three rules to describe the process of growth, invasion, and death. We apply these three generic rules to simulate tumor development—from pixel to cancer—using cellular automata. We then integrate the tumor state into the original computed tomography (CT) images to generate synthetic tumors across different organs. This tumor synthesis approach allows for sampling tumors at multiple stages and analyzing tumor-organ interaction. Clinically, a reader study involving three expert radiologists reveals that the synthetic tumors and their developing trajectories are convincingly realistic. Technically, we generate tumors at varied stages in 9,262 raw, unlabeled CT images sourced from 68 hospitals worldwide. The performance in segmenting tumors in the liver, pancreas, and kidneys exceeds prevailing literature benchmarks, underlining the immense potential of tumor synthesis, especially for earlier cancer detection. The code and models are available at https://github.com/MrGiovanni/Pixel2Cancer
@@ 1 Introduction
Training AI models for image segmentation needs extensive datasets with detailed per-pixel annotations [hatamizadeh2022unetr, isensee2021nnu, zhou2019unet++, zhou2021models, zhou2022interpreting, haghighi2021transferable, feng2020parts2whole]. To reduce manual annotation efforts, data synthesis is a compelling approach for generating enormous synthetic data and annotations for both training [chen2021synthetic, xing2023less, yao2021label, lyu2022pseudo, gao2023synthetic, wang2022anomaly, yu2023good] and evaluating [johnson2017clevr, jordon2022synthetic, liu2023poseexaminer, hu2023synthetic] AI models. This is particularly attractive for segmenting small tumors, which are clinically important to detect at the early stage but rarely occur in public datasets because they are difficult for radiologists to identify.
@@
Preliminary studies have been successful in synthesizing specific disease conditions in medical images, such as synthesizing pulmonary inflammatory lesions in COVID-19 [lyu2022pseudo], retinal diabetic lesions [wang2022anomaly], lung nodules [han2019synthesizing], abdominal tumors [jin2021free, li2023early], and brain tumors [wyatt2022anoddpm]. However, key challenges remain: (i) generating highly realistic tumors, (ii) synthesizing tumors that generalize across multiple organs, and (iii) leveraging synthetic tumors for AI model training. Moreover, no prior work has modeled how tumors develop over time.
@@
In this paper, we revisit a classic technique, cellular automata [neumann1966theory], incorporating novel designs to generate realistic tumors of various shapes. In cellular automata, the basic element is the cell, which our work refers to as a single pixel in the computed tomography (CT) image. We propose Pixel2Cancer, a framework that can generate synthetic tumors guided by three generic rules we have developed [^1] We develop these generic rules by imitating tumor behavior based on medical knowledge (detailed in §2), including the tumor growth and death [wong2015pancreatic, forster2017development, wise2008three, lee2015mathematical], tissue invasion [frieboes2010three, simpson2007simulating, hatzikirou2012go], and interactions with the tumor microenvironments [jessup1993prospects, subramanian2019simulation, tang2014computational].. As a result, Pixel2Cancer can generate multiple stages of tumor development (illustrated in Figure 1), simulate the interaction between tumors and their surrounding environments, exhibit applicability across various organs, and potentially contribute to the estimation of tumor prognosis in longitudinal data.
@@
We have evaluated the Pixel2Cancer from two perspectives. Clinically, reader studies involving three expert radiologists have been conducted, and results have been so convincing that even medical professionals with over ten years of experience can mistake them for real tumors (Table 1). Technically, we have generated synthetic tumors on healthy data to train the segmentation model. The performance in segmenting tumors in the liver, pancreas, and kidneys exceeds previous benchmarks and even model training on real data (Table 3), underlining the immense potential of tumor synthesis, especially for earlier cancer detection/diagnosis (Appendix Figure 4).
@@
This performance is attributable to our key observation: Many tumors, regardless of origin, share underlying growth, invasion, and interaction dynamics. Our generic rules effectively capture these shared processes, enabling tumor synthesis across diverse organs. Benign tumors and early-stage cancers usually grow in a contained manner, pushing against surrounding tissues rather than invading them. In contrast, advanced cancers often exhibit invasive growth, infiltrating surrounding tissues and potentially spreading to distant parts of the body (metastasis) [^2] Using liver cancer as an example, early-stage liver cancer commonly presents as small nodules with a relatively regular shape. In contrast, advanced liver cancer frequently demonstrates infiltration of intrahepatic blood vessels and bile ducts, or distant metastasis. Similar patterns are observed in other solid tumors as well [m2021use].. This suggests the existence of fundamental tumor development principles that transcend organ-specific variations. We hereby contrast Pixel2Cancer with related work to show our innovations and novelties.
@@
    Requiring no manual annotation. Learning-based approaches, such as GAN [goodfellow2014generative] and Diffusion [ho2020denoising, fernandez2023privacy, chen2024towards, du2023boosting], excel in learning tumor representations but require abundant paired tumor data for effective CT image generation. Moreover, the generation process requires extra manual efforts, including the creation of masks to indicate tumor locations and shapes [jin2021free].
@@
    Simulating tumors development. None of the existing synthetic approaches can simulate tumor development, especially in the intricate processes of proliferation and invasion [harpold2007evolution]. The complexity of these processes, influenced by the surrounding environment [tanase2015complexity, wong2015pancreatic], makes it difficult to generate synthetic tumors when interacting with organ tissues and structures, particularly when generalizing to different organs.
@@
    Synthesizing tumors across organs. Modeling-based approaches leverage specialized design and domain expertise to simulate tumor appearances. Although the utilization of modeling-based synthetic tumors eliminates the need for manual annotation [li2023early, hu2022synthetic], they require significant effort for designing proper tumor characteristics customized to a specific organ, therefore limiting generalization, especially across different organs.
@@ 2.2 Simulating Interaction Between Organs and Tumors
Growth and Interaction: As mentioned in §2.1, the tumor will proliferate itself on the tumor population map (self-state $+1$). Increasing populations lead to more space occupation, interacting with and invading surrounding tissues. In the quantified organ, normal tissues (level 1,2,3) can be invaded at different rates. Vessels, boundaries (level 0), and high-density organ tissues (level 4) will impede tumor growth due to the presence of physical barriers. In this scenario, crowded tumor cells (state = 10) exert pressure to invade surrounding tissues, with pressure levels ranging from 0 to n indicating the number of invasion attempts (Figure 2). When the pressure reaches a certain threshold, the tumor can invade these dense tissues. In Figure 2, we also simulate tumor cell death, influenced by factors such as the vascular feeding environment and cell size [rieger2015integrative]. Our model corresponds to cells surrounded by maximal population neighbors (neighbor-states = 10), and certain evolutionary steps (of sufficient size) can lead to cell death.
@@ 2.3 Mapping Tumors to CT Images
After simulating tumor development, the resulting tumor population map needs to be mapped back into CT images (Figure 1). The mapping process is based on a mapping function, eliminating the need for annotations and training processes. It integrates information from the tumor population map and original CT intensity. Tumor population maps guide the mapping process to modify CT intensity values, reflecting different tumor states, including dead cells ($\text{state}=-1$), tumor cells ($\text{state}=10$), normal cells, and blurry boundaries where two cell types mix ($0<\text{state}<10$). To identify the intensity of each tumor pixel, we incorporate typical intensity differences between organ tissues and tumors. Then, we can calculate the intensity by considering the proportion of tumor populations within pixels. To create realistic tumor textures, we introduce 3D Gaussian noise with predefined mean intensity and same standard deviation as organ parenchyma.
3 Experiment & Result
@@ 3.2 Results
Table 1: Visual Turing Test is conducted by three experts, each evaluating 150 CT images, with 50 images allocated for each organ. They were tasked with categorizing each CT image as either real, or synthetic. A lower specificity score indicates a higher number of synthetic tumors being identified as real.
metric
liver
pancreas
kidneys
R1 3-year experience
sensitivity (%)
100
95.0
95.5
specificity (%)
27.3
22.7
26.7
accuracy (%)
60.9
57.1
67.6
R2 7-year experience
sensitivity (%)
94.7
87.5
90.0
specificity (%)
47.8
47.4
56.3
accuracy (%)
69.1
65.7
75.0
R3 10-year experience
sensitivity (%)
100
100
100
specificity (%)
45.4
55.6
57.9
accuracy (%)
68.4
72.4
75.8
  • positives: real tumors ($N$ = 25); negatives: synthetic tumors ($N$ = 25).
Tumor Segmentation Performance: We benchmark Pixel2Cancer against the state-of-the-art modeling-based method [hu2023label] and real-tumor method [liu2023clip]. Table 2 demonstrates Pixel2Cancer’s superiority in liver segmentation with DSC of $58.9\%$, and NSD of $63.7\%$. We also achieve superior performance with real tumors in the pancreas and kidney, with DSCs of $60.9\%$ and $73.2\%$, respectively. Compared to models trained on real liver tumors, Pixel2Cancer outperforms by $5.7\%$ in NSD. We also surpass Hu et al. by $4.4\%$ in DSC and $6.1\%$ in NSD.
organ tumor U-Net Swin UNETR nnU-Net
DSC/NSD
SD/HD
DSC/NSD
SD/HD
DSC/NSD
SD/HD
(%) $\uparrow$
(mm) $\downarrow$
(%) $\uparrow$
(mm) $\downarrow$
(%) $\uparrow$
(mm) $\downarrow$
liver
real tumors
56.7/58.0
23.2/61.4
53.5/55.9
21.3/57.8
56.2/55.3
24.6/58.3
Hu et al.
54.5/57.6
23.8/58.8
52.3/56.5
22.9/56.9
53.7/56.1
22.5/57.2
Pixel2Cancer
58.9/63.7 17.9/52.4 56.7/62.5 18.7/51.3 57.9/63.2 18.9/52.7
pancreas
real tumors
57.8/56.5
13.1/47.7
56.7/52.8
24.6/53.9
56.8/52.1
14.5/44.6
Hu et al.
54.1/52.2
15.7/49.3
53.6/54.9
22.5/47.4
54.6/52.4
17.1/48.0
Pixel2Cancer
60.9/57.1 12.4/43.5 59.3/59.5 20.4/40.7 59.8/56.9 13.3/41.4
kidney
real tumors
71.3/62.8
27.2/64.3
70.7/61.2
19.8/57.1
65.2/58.1
25.6/59.3
Hu et al.
63.2/55.4
35.1/69.0
61.7/52.3
26.2/61.6
55.5/49.9
27.9/62.7
Pixel2Cancer
73.2/65.0 13.6/40.9 73.9/63.5 15.9/45.7 67.6/60.1 14.8/42.2
  • DSC - dice similarity coefficient; NSD - normalized surface dice.
  • SD - surface distance; HD - Hausdorff distance.
@@
Figure 3: (a) Different Condition: We evaluate the influence of different tumor sizes, intensities, and roundness on liver tumors to assess the robustness of Pixel2Cancer. (b) Early Tumor Detection: We report Sensitivity of tumor detection among real tumors [liu2023clip], Hu et al. [hu2023label], and Pixel2Cancer in three organs. (c) Effectiveness of Rules: We evaluated the performance of each generic rule on liver tumors, demonstrating the effectiveness of our rule design in tumor simulation.
Ability in Small Tumor Detection: Early detection of small tumors is essential for timely cancer diagnosis, but real datasets often lack sufficient instances due to the lack of CT scans in the early stages. Figure 3(b) and Appendix Figure 4 illustrates the performance of small tumor detection. Training solely on our synthetic tumor outperforms models trained on real tumors [liu2023clip] and Hu et al. [hu2023label], highlighting the potential of synthetic tumors for improving early cancer detection.
Ablation Studies: (1) We evaluated how different tumor conditions affect the segmentation results. Pixel2Cancer exhibits greater variability and robustness across different conditions compared to real tumors (Figure 3(a)). (2) We also evaluated the effectiveness of generic rules on liver tumors (Figure 3(c)), basic growth rules yielded $52.1\%$ NSD. Incorporating interaction and invasion achieved performance comparable to that of real data. Adding the tumor cell death rule outperforms real tumors by $4.9\%$ in NSD, demonstrating the effectiveness of each generic rule.
@@ 4 Conclusion & Discussion
In this paper, we highlight the potential of Pixel2Cancer for tumor synthesis, data augmentation, small tumor detection, and accurate boundary segmentation. Despite simulating tumor behaviors and interactions with organ tissues, including the bending of vessels and the deformation of organ boundaries, there are still various mechanisms of accompanying changes in organs caused by tumors that we have not yet considered, including focal pancreas parenchymal atrophy [miura2021focal], pancreatic duct [xia2022felix], hepatic capsular retraction [tan2016causes], and splenomegaly. In future work, we plan to devise organ rules within our approach to accurately simulate changes induced by tumors when generating synthetic tumors.

### v2

@@ Abstract
AI for cancer detection encounters the bottleneck of data scarcity, annotation difficulty, and low prevalence of early tumors. Tumor synthesis seeks to create artificial tumors in medical images, which can greatly diversify the data and annotations for AI training. However, current tumor synthesis approaches are not applicable across different organs due to their need for specific expertise and design. This paper establishes a set of generic rules to simulate tumor development. Each cell (pixel) is initially assigned a state between zero and ten to represent the tumor population, and a tumor can be developed based on three rules to describe the process of growth, invasion, and death. We apply these three generic rules to simulate tumor development—from pixel to cancer—using cellular automata. We then integrate the tumor state into the original computed tomography (CT) images to generate synthetic tumors across different organs. This tumor synthesis approach allows for sampling tumors at multiple stages and analyzing tumor-organ interaction. Clinically, a reader study involving three expert radiologists reveals that the synthetic tumors and their developing trajectories are convincingly realistic. Technically, we analyze and simulate tumor development at various stages using 9,262 raw, unlabeled CT images sourced from 68 hospitals worldwide. The performance in segmenting tumors in the liver, pancreas, and kidneys exceeds prevailing literature benchmarks, underlining the immense potential of tumor synthesis, especially for earlier cancer detection. The code and models are available at https://github.com/MrGiovanni/Pixel2Cancer
@@ 1 Introduction
Training AI models for image segmentation needs extensive datasets with detailed per-pixel annotations To reduce manual annotation efforts, data synthesis is a compelling approach for generating enormous synthetic data and annotations for both training [lyu2022pseudo, hu2023label] and evaluating [hu2023synthetic] AI models. This is particularly attractive for segmenting small tumors, which are clinically important to detect at the early stage but rarely occur in public datasets because they are difficult for radiologists to identify.
@@
Preliminary studies have been successful in synthesizing specific disease conditions in medical images, such as synthesizing pulmonary inflammatory lesions in COVID-19 [lyu2022pseudo], lung nodules [han2019synthesizing], abdominal tumors [jin2021free, chen2024towards], and brain tumors [wyatt2022anoddpm]. However, key challenges remain: (i) generating highly realistic tumors, (ii) synthesizing tumors that generalize across multiple organs, and (iii) leveraging synthetic tumors for AI model training. Moreover, while some mathematical studies [forster2017development, menze2011generative] have modeled tumor development in simplified hypothetical environments, they cannot generate realistic synthetic data for model training.
@@
In this paper, we revisit a classic technique, cellular automata [wolfram1983statistical], incorporating novel designs to generate realistic tumors of various shapes. In cellular automata, the basic element is the cell, which our work refers to as a single pixel in the computed tomography (CT) image. We propose Pixel2Cancer, a framework that can generate synthetic tumors guided by three generic rules we have developed [^1] We develop these generic rules by imitating tumor behavior based on medical knowledge (detailed in §2), including tumor growth and death [forster2017development], tissue invasion [frieboes2010three], and interactions with the tumor microenvironments [subramanian2019simulation].. As a result, Pixel2Cancer can generate multiple stages of tumor development (illustrated in Figure 1), simulate the interaction between tumors and their surrounding environments, exhibit applicability across various organs, and potentially contribute to the estimation of tumor prognosis in longitudinal data.
@@
We have evaluated the Pixel2Cancer from two perspectives. Clinically, reader studies involving three expert radiologists have been conducted, and results have been so convincing that even medical professionals with over ten years of experience can mistake them for real tumors (Table 1). Technically, we have generated synthetic tumors on healthy data to train the segmentation model. The performance in segmenting tumors in the liver, pancreas, and kidneys exceeds previous benchmarks and even model training on real data (Appendix Table 3), underlining the immense potential of tumor synthesis, especially for earlier cancer detection/diagnosis (Appendix Figure 4).
@@
This performance is attributable to our key observation: Many tumors, regardless of origin, share underlying growth, invasion, and interaction dynamics. Our generic rules effectively capture these shared processes, enabling tumor synthesis across diverse organs. Benign tumors and early-stage cancers usually grow in a contained manner, pushing against surrounding tissues rather than invading them. In contrast, advanced cancers often exhibit invasive growth, infiltrating surrounding tissues and potentially spreading to distant parts of the body (metastasis) [^2] Using liver cancer as an example, early-stage liver cancer commonly presents as small nodules with a relatively regular shape. In contrast, advanced liver cancer frequently demonstrates infiltration of intrahepatic blood vessels and bile ducts, or distant metastasis. Similar patterns are observed in other solid tumors as well [m2021use].. Furthermore, we found that the 9,262 raw CT images we used were captured during the venous phase, during which a significant proportion of tumors appeared hypo-intense (dark). Therefore, we chose to simulate this typical presentation, closely aligning with clinical practice. This observation implies the existence of fundamental tumor development principles that transcend organ-specific variations. We hereby contrast Pixel2Cancer with related work to show our innovations and novelties.
@@
    Requiring no manual annotation. Learning-based approaches, such as GAN [goodfellow2014generative] and Diffusion [ho2020denoising], excel in learning tumor representations but require abundant paired tumor data for effective CT image generation. Moreover, the generation process requires extra manual efforts, including the creation of masks to indicate tumor locations and shapes [jin2021free].
@@
    Simulating tumors development. None of the existing synthetic approaches can simulate tumor development, especially in the intricate processes of proliferation and invasion [harpold2007evolution]. The complexity of these processes, influenced by the surrounding environment [tanase2015complexity], makes it difficult to generate synthetic tumors when interacting with organ tissues and structures, particularly when generalizing to different organs.
@@
    Synthesizing tumors across organs. Modeling-based approaches leverage specialized design and domain expertise to simulate tumor appearances. Although the utilization of modeling-based synthetic tumors eliminates the need for manual annotation [hu2023label], they require significant effort for designing proper tumor characteristics customized to a specific organ, therefore limiting generalization, especially across different organs.
@@ 2.2 Simulating Interaction Between Organs and Tumors
Growth and Interaction: As mentioned in §2.1, the tumor will proliferate itself on the tumor population map (self-state $+1$). Increasing populations lead to greater space occupation, resulting in interactions with and invasion of surrounding tissues. Crowded tumor cells will randomly choose directions and attempt to invade. When interacting with organs, tumors can directly invade normal tissues (level 1,2,3) at different rates. However, the presence of physical barriers such as vessels, boundaries (level 0), and high-density organ tissues (level 4) can impede tumor growth. In this case, crowded tumor cells (state = 10) exert pressure as they attempt to invade surrounding tissues. The pressure levels range from 0 to n, indicating the number of invasion attempts (Figure 2). When the pressure of the target direction reaches the threshold, tumor cells can invade these high-density tissues. In Figure 2, we also simulate tumor cell death, influenced by factors such as the vascular feeding environment and cell size [rieger2015integrative]. Our model corresponds to cells surrounded by maximal population neighbors (neighbor-states = 10), and certain evolutionary steps (of sufficient size) can lead to cell death.
@@ 2.3 Mapping Tumors to CT Images
After simulating tumor development, the resulting tumor population map needs to be mapped back into CT images (Figure 1). The mapping process is based on a mapping function, eliminating the need for annotations and training processes. It integrates information from the tumor population map and original CT intensity. Tumor population maps guide the mapping process to modify CT intensity values, reflecting different tumor states, including dead cells ($\text{state}=-1$), tumor cells ($\text{state}=10$), normal cells, and blurry boundaries where two cell types mix ($0<\text{state}<10$). To identify the intensity of each tumor pixel, we incorporate typical intensity differences between organ tissues and tumors. Then, we can calculate the intensity by considering the proportion of tumor populations within pixels. To create realistic tumor textures, we introduce 3D Gaussian noise with predefined mean intensity and standard deviation as organ parenchyma.
Table 1: Visual Turing Test involved three experts each evaluating 150 CT images, with 50 per organ. They were tasked with categorizing each CT image as either real, or synthetic. A lower specificity score indicates a higher number of synthetic tumors being identified as real. We also provide typical examples of synthetic tumors recognized correctly and incorrectly. The red dot indicates the human expert believes this is a synthetic tumor, while the green dot indicates they believe it is a real tumor.
metric liver pancreas kidneys
R1 3-year experience sensitivity (%) 100 95.0 95.5
specificity (%) 27.3 22.7 26.7
accuracy (%) 60.9 57.1 67.6
R2 7-year experience sensitivity (%) 94.7 87.5 90.0
specificity (%) 47.8 47.4 56.3
accuracy (%) 69.1 65.7 75.0
R3 10-year experience sensitivity (%) 100 100 100
specificity (%) 45.4 55.6 57.9
accuracy (%) 68.4 72.4 75.8
  • positives: real tumors ($N$ = 25); negatives: synthetic tumors ($N$ = 25).
3 Fake Tumors, Real Results
@@ 3.1 Experiment
Hyper-parameter in Pixel2Cancer: We hereby specify the hyper-parameters when generalizing Pixel2Cancer to new organs. The quantification process requires setting the typical highest and lowest HU values for organs. The growth and interaction process requires setting the invasion pressure threshold based on organ softness and the maximum growth steps based on organ size. The mapping function requires setting the typical tumor HU values for different organs.
organ tumor U-Net Swin UNETR nnU-Net
DSC/NSD SD/HD DSC/NSD SD/HD DSC/NSD SD/HD
(%) $\uparrow$ (mm) $\downarrow$ (%) $\uparrow$ (mm) $\downarrow$ (%) $\uparrow$ (mm) $\downarrow$
liver real tumors 56.7/58.0 23.2/61.4 53.5/55.9 21.3/57.8 56.2/55.3 24.6/58.3
Hu et al. 54.5/57.6 23.8/58.8 52.3/56.5 22.9/56.9 53.7/56.1 22.5/57.2
Pixel2Cancer 58.9/63.7 17.9/52.4 56.7/62.5 18.7/51.3 57.9/63.2 18.9/52.7
pancreas real tumors 57.8/56.5 13.1/47.7 56.7/52.8 24.6/53.9 56.8/52.1 14.5/44.6
Hu et al. 54.1/52.2 15.7/49.3 53.6/54.9 22.5/47.4 54.6/52.4 17.1/48.0
Pixel2Cancer 60.9/57.1 12.4/43.5 59.3/59.5 20.4/40.7 59.8/56.9 13.3/41.4
kidney real tumors 71.3/62.8 27.2/64.3 70.7/61.2 19.8/57.1 65.2/58.1 25.6/59.3
Hu et al. 63.2/55.4 35.1/69.0 61.7/52.3 26.2/61.6 55.5/49.9 27.9/62.7
Pixel2Cancer 73.2/65.0 13.6/40.9 73.9/63.5 15.9/45.7 67.6/60.1 14.8/42.2
  • DSC - dice similarity coefficient; NSD - normalized surface dice.
  • SD - surface distance; HD - Hausdorff distance.
@@ 3.2 Results
Tumor Segmentation Performance: We benchmark Pixel2Cancer against the state-of-the-art modeling-based method [hu2023label] and real-tumor method. Table 2 demonstrates Pixel2Cancer’s superiority in liver segmentation with DSC of $58.9\%$, and NSD of $63.7\%$. We also achieve superior performance with real tumors in the pancreas and kidney, with DSCs of $60.9\%$ and $73.2\%$, respectively. Compared to models trained on real liver tumors, Pixel2Cancer outperforms by $5.7\%$ in NSD. We also surpass Hu et al. by $4.4\%$ in DSC and $6.1\%$ in NSD.
@@
Figure 3: (a) Different condition: We evaluate the influence of different tumor sizes, intensities, and roundness on liver tumors to assess the robustness of Pixel2Cancer. (b) Early tumor detection: We report the Sensitivity of U-Net tumor detection among real tumors, Hu et al. [hu2023label], and Pixel2Cancer in three organs. (c) Effectiveness of rules: We evaluated the performance of each generic rule on liver tumors, demonstrating the effectiveness of our rule design in tumor simulation.
Ability in Small Tumor Detection: Early detection of small tumors is essential for timely cancer diagnosis, but real datasets often lack sufficient instances due to the lack of CT scans in the early stages. Figure 3(b) and Appendix Figure 4 illustrates the performance of small tumor detection. Training solely on our synthetic tumor outperforms models trained on real tumors and Hu et al. [hu2023label], highlighting the potential of synthetic tumors for early cancer detection.
Ablation Studies: (1) We evaluated the impact of various tumor conditions on the performance of the model. Pixel2Cancer exhibits greater variability and robustness across different conditions compared to real tumors (Figure 3(a)). (2) We evaluated the effectiveness of generic rules on liver tumors (Figure 3 (c)) Basic growth rules yielded $52.1\%$ NSD. Incorporating interaction and invasion matched real data performance. Adding the tumor death rule outperformed real tumors by $4.9\%$ in NSD, demonstrating the effectiveness of these generic rules.
@@ 4 Conclusion & Discussion
In this paper, we highlight the potential of Pixel2Cancer for tumor synthesis, data augmentation, small tumor detection, and accurate boundary segmentation. Despite simulating tumor behaviors and interactions with organ tissues, including the bending of vessels and the deformation of organ boundaries, there are still various mechanisms of accompanying changes in organs caused by tumors that we have not yet considered, including focal pancreas parenchymal atrophy hepatic capsular retraction [tan2016causes], and splenomegaly. In future work, we plan to devise organ rules within our approach to accurately simulate changes induced by tumors when generating synthetic tumors.
{credits}